How much water must be added to 16 liters of a 30% solution to get a solution that is 22% acid? Please show the steps, so that I may do the same with a similar problem.
Thank you.
Byron

Answer by Alan3354(69443)

How much water must be added to 16 liters of a 30% solution to get a solution that is 22% acid?
----------------
16 liters of 30% = 4.8 liters of acid.
Add water until that 4.8 liters is 22% of the total.
4.8:22::x:100
22x = 480
x = 240/11 liters = the total volume of 22%
240/11 - 16 = 64/11 liters of water to add
64/11 = ~ 5.818 liters of water
-----------------
Note: If you add 16 liters of water, the %age would be 1/2, or 15%, so the answer looks good.